# OptMerge: Unifying Multimodal LLM Capabilities and Modalities via Model Merging

### v2

@@ Abstract
Foundation models update slowly due to resource-intensive training, whereas domain-specific models evolve rapidly between releases. Model merging seeks to combine multiple expert models into a single, more capable model, reducing storage and serving costs while supporting decentralized development. Despite its potential, previous studies have primarily focused on merging visual classification models or Large Language Models (LLMs) for code and math tasks. Recently, Multimodal LLMs (MLLMs) that extend LLMs through large-scale multimodal training have gained traction. However, there lacks a benchmark for model merging research that clearly divides the tasks for MLLM training and evaluation. In this paper, (i) we introduce a model merging benchmark for MLLMs, which includes multiple tasks such as VQA, Geometry, Chart, OCR, and Grounding, studying both LoRA and full fine-tuning models. Moreover, we explore how model merging can combine different modalities (e.g., vision-language, audio-language, and video-language models), moving toward the Omni-language model. (ii) We implement 10 model merging algorithms on the benchmark. Furthermore, we propose a novel method that removes noise from task vectors and robustly optimizes the merged vector based on a loss defined over task vector interactions, achieving an average performance gain of 2.48%. (iii) We find that model merging offers a promising way for building improved MLLMs without requiring training data. Our results also demonstrate that the complementarity among multiple modalities outperforms individual modalities. All code and checkpoints are publicly available here.
@@ 1 Introduction
Recently, Multimodal Large Language Models (MLLMs) that extend LLMs with broader capabilities through large-scale multimodal training have gained traction. Model merging offers a cost-effective way to combine fine-tuned MLLMs with task-specific skills into a unified model (see Fig.˜1). However, there lacks a benchmark for model merging research that clearly divides the tasks for MLLM training and evaluation. Specifically, AdaMMS (Du et al., 2025) proposes an unsupervised hyper-parameter selection method, but can only merge two MLLMs at a time. For example, merging LLaVA-OneVision-Qwen (Li et al., 2025a) into Qwen2-VL (Wang et al., 2024b) on the Qwen2 architecture. UQ-Merge (Qu et al., 2025) treats each fine-tuning dataset in LLaVA-v1.5 (Liu et al., 2024a) as a separate task without categorization of MLLM capabilities, fine-tuning the base model for each dataset and using LLaVA-v1.5 as the mixture training baseline.
In this paper, we introduce a model merging benchmark for MLLMs, which includes diverse specialized models such as VQA, Geometry, Chart, OCR, and Grounding. For each corresponding task, we collect comprehensive public datasets with at least 100k samples to ensure effective supervised fine-tuning, and select corresponding benchmarks to evaluate distinct capabilities. We derive an upper bound on the error between the merged model and expert models, proving that merging performance is influenced by the learning rate and iterations, which control the extent of parameter drift. Smaller parameter changes sacrifice task performance but lead to easier merging. We choose two types of vision-language models: InternVL2.5 and Qwen2-VL, providing both LoRA and full fine-tuning checkpoints. Moreover, most existing MLLMs specialize in dual modalities, and incorporating new modality encoders requires re-training on new modality-text data. Generating high-quality multimodal instruction data is resource-consuming (Jiang et al., 2024). Therefore, we explore how model merging can efficiently combine different modalities (e.g., vision-language, audio-language, and video-language models), moving toward the Omni-language model. This offers a data-free way to reuse and integrate modality-specific encoders into a unified LLM.
@@
In summary, our main contributions are three-fold:
@@ 2 Related Work
Static merging can be further divided into data-free methods and test-time adaptation methods. Data-free methods merge fine-tuned models without requiring additional data. We categorize these methods into four groups: (i) Linear interpolation methods that perform arithmetic operations on task vectors (Wortsman et al., 2022; Ilharco et al., 2023; Goddard et al., 2024; Chen et al., 2025); (ii) Sparsification-based methods that reduce redundancy in task vectors (Yadav et al., 2023; Yu et al., 2024; He et al., 2025); (iii) SVD-based methods that identify and exploit the low-rank features of task vectors (Gargiulo et al., 2025; Marczak et al., 2025; Choi et al., 2024; Stoica et al., 2025); and (iv) Optimization-based methods that optimize task vectors via gradient descent (Wei et al., 2025b; Cheng et al., 2025). Test-time adaptation (Yang et al., 2024d; c; Daheim et al., 2024) assumes access to unlabeled test datasets, which can be considered a form of transductive learning.
@@
Several approaches similar to ours aim to merge multiple MLLMs to improve multi-task performance. AdaMMS (Du et al., 2025) proposes an unsupervised hyperparameter selection method for model merging. However, it requires generating responses for each candidate hyperparameter, making it time-consuming and assuming test set availability during merging. Furthermore, it can only merge two models at a time. For example, merging LLaVA-OneVision-Qwen into Qwen2-VL on the Qwen2 architecture, or merging LLaVA-v1.5 into CogVLM-Chat on the LLaMA architecture. UQ-Merge (Qu et al., 2025) considers uncertainty quantification on text and vision inputs to examine MLLM prediction confidence, requiring unlabeled test sets to calculate prediction and determine merging sequence. This approach measures uncertainty across all candidate models and repeatedly evaluates merged models to find optimal combinations. UQ-Merge treats each fine-tuning dataset in LLaVA-v1.5 (Liu et al., 2024a) as a separate task without categorization of MLLM capabilities, fine-tuning the base model for each dataset and using LLaVA-v1.5 as the mixture training baseline. In contrast, our benchmark collects more comprehensive data with clearer MLLM task divisions for fine-tuning, and we propose a data-free method that requires no hyperparameter search.
@@ 3.1 Merging Baselines
Linear interpolation methods: Weight Averaging (Wortsman et al., 2022) simply averages the weights of models fine-tuned on different tasks. Task Arithmetic (Ilharco et al., 2023) computes task vectors $\bm{\tau}_{i}=\bm{\theta}_{i}-\bm{\theta}_{0}$ for individual tasks and sums them to form a multi-task vector $\bm{\tau}_{m}=\sum_{i=1}^{n}\bm{\tau}_{i}$. This vector is scaled by a coefficient $\lambda$ and added to the base model $\bm{\theta}_{0}$ to obtain the merged model.
@@ 3.2 Parameter Changes during Fine-tuning Matter
Model merging exhibits sensitivity to task vectors $\bm{\tau}_{i}$ (i.e., parameter changes between fine-tuned models and the base model). Several studies (Yu et al., 2024; Li et al., 2025b) demonstrate that less intensive fine-tuning can yield superior merging performance, even when these models achieve lower accuracy on their respective tasks. In App. B.1, we conduct experiments on the impact of fine-tuning steps on merging performance, and observe that performance tends to rise initially and then decline. This counter-intuitive finding suggests that higher-performing expert models do not necessarily produce better merging outcomes. Fine-tuned models tend to converge around the base model in parameter space (Merlin et al., 2023; Chung et al., 2024). When constructing our benchmark, we reduce parameter changes by lowering the learning rate while ensuring performance improvements on specific tasks. We analyze the upper bound of the loss incurred by model merging:
@@
Let $\bm{\theta}_{0}$ be the initial parameter. For each task $i$, after $T$ gradient steps with constant learning rate $\eta$, we denote the task vector as $\bm{\tau}_{i}=\bm{\theta}_{i}-\bm{\theta}_{0}$. Considering that most methods can be viewed as extensions of linear combinations of task vectors, let $\bm{\tau}_{m}=\sum_{j}\alpha_{j}\bm{\tau}_{j}$ denote the merged vector. The loss on task $i$ is denoted by $\mathcal{L}_{i}(\Theta)$, which is $\mathcal{C}_{i}$-Lipschitz continuous. Assuming a constraint on the second moment of the gradient, then:
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\mathcal{O}(\eta T)$ (2)
This indicates that both the learning rate and iterations influence model merging results. Please refer to App. A for detailed assumptions and proofs.
Remark. Theorem˜3.1 provides the first theoretical explanation of how model fine-tuning affects merging performance. It supports previous empirical observations: fine-tuned models in current benchmarks typically remain within the same basin near the base model. For example, merging Qwen2.5-Math (Yang et al., 2024a) and Qwen2.5-Coder (Hui et al., 2024) yields poor performance. This is likely due to excessive post-training, which causes significant parameter drift. This insight suggests that poor merging results may not reflect algorithmic flaws, but rather issues with the fine-tuned models. When selecting models from Hugging Face, it’s helpful to choose fine-tuned models that are better suited for merging to minimize multi-task degradation.
@@
We examine the weight magnitude distribution of task vectors across our benchmark (see Fig.˜2 (a-b)). Our analysis reveals that InternVL2.5, which undergoes full fine-tuning, exhibits a right-skewed distribution. In contrast, Qwen2-VL, fine-tuned using LoRA, displays a multi-modal distribution. This is due to the low-rank nature and scaling factors of LoRA, which constrain the task vectors to be linear combinations in a reduced subspace, causing them to cluster along a few dominant magnitudes. Both models demonstrate distinct magnitude distribution patterns across different tasks. We also compute the normalized Frobenius norm of parameters (i.e., divided by the number of parameters). As shown in Fig.˜2 (c-d), the Frobenius norm varies significantly across tasks and layers, which presents challenges that we will address in our approach. The small task vector magnitudes suggest that fine-tuned models and base models exist in adjacent regions of the loss landscape with linear connectivity (Wu et al., 2023), facilitating effective model merging (Ortiz-Jimenez et al., 2023).
@@ 4.1 Merging Full Fine-tuning Models
By truncating singular values, we preserve critical features $V_{1:k}^{\top}$, which is similar to selecting principal components in Principal Components Analysis (PCA) (Abdi & Williams, 2010). This yields more accurate estimates of $\bm{x}{i,l}$ than using $(\bm{\tau}_{i,l})^{\top}$.
@@ 5.1 Benchmark Details
VQA 588K GQA (en) (Hudson & Manning, 2019), VQAv2 (en) (Goyal et al., 2017), OKVQA (en) (Marino et al., 2019), LLaVA-Instruct (zh) (Liu et al., 2024a), CogVLM-Singleround (en&zh) (Wang et al., 2024c), CogVLM-Multiround (en&zh) (Wang et al., 2024c)
Geometry 190K GeoQA+ (zh) (Cao & Xiao, 2022), G-LLaVA (en) (Gao et al., 2023)
@@
Merging details. Following Task Arithmetic (Ilharco et al., 2023), we employ a single coefficient $\lambda$ to scale the merged vector before adding it to the base model. For all model merging methods, we determine the optimal merging coefficient $\lambda$ by searching within the range [0.1, 0.3, 0.5, 0.7, 1.0, 1.5]. In our implementation, the rank size $k$ in Eq.˜4 is simply defined as the rank of each task vector divided by the number of tasks (i.e., 5). We use the Adam optimizer with a learning rate of 1e-5 for InternVL while applying the SGD optimizer with a learning rate of 1e-4 for QwenVL. The number of optimization iterations is set to 300. We apply our method exclusively to the linear layer in the model.
@@ 5.2 Experimental Results
Table 2: Capability merging results on InternVL2.5 (full fine-tuning) across multiple tasks. For the merging methods, we highlight the best score in bold and the second-best score with underlining.
@@
Table 3: Capability merging results on Qwen2-VL (LoRA fine-tuning) across multiple tasks. For the merging methods, we highlight the best score in bold and the second-best score with underlining.
@@
Improved task vector optimization. Our method enhances task vector optimization stability, achieving optimal results. In Table˜4, we evaluate each component’s contribution to overall performance. Starting with WUDI Merging, we incrementally add one component at a time, reporting performance for both LoRA model merging (Qwen2-VL) and modality merging (Vicuna-7B). Replacing Adam with SGD alone does not necessarily improve performance; however, when combined with initializing the merged vector using the mean of task vectors, we observe a significant 4.43% improvement. Low-rank approximation further enhances performance, demonstrating its effectiveness in preserving critical knowledge from task vectors while maintaining the stability of the Frobenius norm. For full fine-tuned models, Tables˜2 and 6 show average improvements of 0.44% and 1.9% for OptMerge over WUDI Merging, respectively. This highlights the necessity of Eq.˜4 over Eq.˜1 for denoising task vectors and achieving robust merged-vector optimization.

### v3

@@ Abstract
Foundation models update slowly due to resource-intensive training, whereas domain-specific models evolve rapidly between releases. Model merging seeks to combine multiple expert models into a single, more capable model, reducing storage and serving costs while supporting decentralized development. Despite its potential, previous studies have primarily focused on merging visual classification models or Large Language Models (LLMs) for code and math tasks. Recently, Multimodal LLMs (MLLMs) that extend LLMs through large-scale multimodal training have gained traction. However, no benchmark exists for model merging research that clearly divides the tasks of MLLM training and evaluation. In this paper, (i) we introduce a model merging benchmark for MLLMs, which includes multiple tasks such as VQA, Geometry, Chart, OCR, and Grounding, studying both LoRA and full fine-tuning models. Moreover, we explore how model merging can combine different modalities (e.g., vision-language, audio-language, and video-language models), moving toward the Omni-language model. (ii) We implement 10 model merging algorithms on the benchmark. Furthermore, we propose a novel method that removes noise from task vectors and robustly optimizes the merged vector based on a loss defined over task vector interactions, achieving an average performance gain of 2.48%. (iii) We find that model merging offers a promising way for building improved MLLMs without requiring training data. Our results also demonstrate that the complementarity among multiple modalities outperforms individual modalities. All code and checkpoints are publicly available here.
@@ 1 Introduction
Recently, Multimodal Large Language Models (MLLMs) that extend LLMs with broader capabilities through large-scale multimodal training have gained traction. Model merging offers a cost-effective way to combine fine-tuned MLLMs with task-specific skills into a unified model (see Fig.˜1). However, no benchmark exists for model merging research that clearly divides the tasks of MLLM training and evaluation. Specifically, AdaMMS (Du et al., 2025b) proposes an unsupervised hyper-parameter selection method, but can only merge two MLLMs at a time. For example, merging LLaVA-OneVision-Qwen (Li et al., 2025a) into Qwen2-VL (Wang et al., 2024b) on the Qwen2 architecture. UQ-Merge (Qu et al., 2025) treats each fine-tuning dataset in LLaVA-v1.5 (Liu et al., 2024a) as a separate task without categorization of MLLM capabilities, fine-tuning the base model for each dataset, and using LLaVA-v1.5 as the mixture training baseline.
In this paper, we introduce a model merging benchmark for MLLMs, which includes diverse specialized models such as VQA, Geometry, Chart, OCR, and Grounding. For each corresponding task, we collect comprehensive public datasets with at least 100k samples to ensure effective supervised fine-tuning, and select corresponding benchmarks to evaluate distinct capabilities. We derive an upper bound on the error between the merged model and expert models, proving that merging performance is influenced by the learning rate and iterations, which control the extent of parameter drift. Smaller parameter changes sacrifice task performance but lead to more effortless merging. We choose two types of vision-language models: InternVL2.5 and Qwen2-VL, providing both LoRA and full fine-tuning checkpoints. Moreover, most existing MLLMs specialize in dual modalities, and incorporating new modality encoders requires re-training on new modality-text data. Generating high-quality multimodal instruction data is resource-consuming (Jiang et al., 2024). Therefore, we explore how model merging can efficiently combine different modalities (e.g., vision-language, audio-language, and video-language models), moving toward the Omni-language model. This offers a data-free way to reuse and integrate modality-specific encoders into a unified LLM.
@@
In summary, our main contributions are threefold:
@@ 2 Related Work
Static merging can be further divided into data-free methods and test-time adaptation methods. Data-free methods merge fine-tuned models without requiring additional data. We categorize these methods into four groups: (i) Linear interpolation methods that perform arithmetic operations on task vectors (Wortsman et al., 2022; Ilharco et al., 2023; Goddard et al., 2024; Chen et al., 2025); (ii) Sparsification-based methods that reduce redundancy in task vectors (Yadav et al., 2023; Yu et al., 2024; He et al., 2025); (iii) SVD-based methods that identify and exploit the low-rank features of task vectors (Gargiulo et al., 2025; Marczak et al., 2025; Choi et al., 2024; Stoica et al., 2025); and (iv) Optimization-based methods that optimize task vectors via gradient descent (Wei et al., 2025c; Cheng et al., 2025). Test-time adaptation (Yang et al., 2024d; c; Daheim et al., 2024) assumes access to unlabeled test datasets, which can be considered a form of transductive learning.
@@
Several approaches similar to ours aim to merge multiple MLLMs to improve multi-task performance. AdaMMS (Du et al., 2025b) proposes an unsupervised hyperparameter selection method for model merging. However, it requires generating responses for each candidate hyperparameter, making it time-consuming and assuming test set availability during merging. Furthermore, it can only merge two models at a time. For example, merging LLaVA-OneVision-Qwen into Qwen2-VL on the Qwen2 architecture, or merging LLaVA-v1.5 into CogVLM-Chat on the LLaMA architecture. UQ-Merge (Qu et al., 2025) considers uncertainty quantification on text and vision inputs to examine MLLM prediction confidence, requiring unlabeled test sets to calculate prediction and determine merging sequence. This approach measures uncertainty across all candidate models and repeatedly evaluates merged models to find optimal combinations. UQ-Merge treats each fine-tuning dataset in LLaVA-v1.5 (Liu et al., 2024a) as a separate task without categorization of MLLM capabilities, fine-tuning the base model for each dataset and using LLaVA-v1.5 as the mixture training baseline. In contrast, our benchmark collects more comprehensive data with clearer MLLM task divisions for fine-tuning, and we propose a data-free method that requires no hyperparameter search.
@@ 3.1 Merging Baselines
Linear interpolation methods: Weight Averaging (Wortsman et al., 2022) averages the weights of models fine-tuned on different tasks. Task Arithmetic (Ilharco et al., 2023) computes task vectors $\bm{\tau}_{i}=\bm{\theta}_{i}-\bm{\theta}_{0}$ for individual tasks and sums them to form a multi-task vector $\bm{\tau}_{m}=\sum_{i=1}^{n}\bm{\tau}_{i}$. This vector is scaled by a coefficient $\lambda$ and added to the base model $\bm{\theta}_{0}$ to obtain the merged model.
@@ 3.2 Parameter Changes during Fine-tuning Matter
Model merging exhibits sensitivity to task vectors $\bm{\tau}_{i}$ (i.e., parameter changes between fine-tuned models and the base model). Several studies (Yu et al., 2024; Li et al., 2025b) demonstrate that less intensive fine-tuning can yield superior merging performance, even when these models achieve lower accuracy on their respective tasks. In App. B.1, we conduct experiments on the impact of fine-tuning steps on merging performance, and observe that performance tends to rise initially and then decline. This counterintuitive finding suggests that higher-performing expert models do not necessarily produce better merging outcomes. Fine-tuned models tend to converge around the base model in parameter space (Merlin et al., 2023; Chung et al., 2024). When constructing our benchmark, we minimize parameter changes by adjusting the learning rate while maintaining performance improvements on specific tasks. We analyze the upper bound of the loss incurred by model merging:
@@
Consider task $i$ trained for $T$ iterations of gradient descent with a fixed step size $\eta\in(0,1/L]$, where $L$ is the Lipschitz constant. Let $\gamma:=1-\eta\mu\in(0,1)$ denote the PL convergence factor. Then the merged update $\bm{\tau}_{m}:=\sum_{j=1}^{m}\alpha_{j}\bm{\tau}_{j}$ satisfies
$\mathcal{L}_{i}(\bm{\Theta}+\bm{\tau}_{m})\leq C_{i}+\mathcal{O}(\gamma^{T})+\mathcal{O}(\delta\,\eta T)+\mathcal{O}(\eta^{2}T^{2}),$
where $\mathcal{O}(\gamma^{T})$ is the residual error from incomplete convergence on task $i$, $\mathcal{O}(\delta\,\eta T)$ is the cross-task interference term, and $\mathcal{O}(\eta^{2}T^{2})$ is the curvature term from $L$-smoothness. This indicates that both the learning rate and iterations influence model merging results. Please refer to App. A for detailed assumptions and proofs.
Remark. Theorem˜3.1 provides the first theoretical explanation of how model fine-tuning affects merging performance. The target task’s gains dominate in the early training stage, but as convergence approaches, cross-task interference and curvature errors (growing with $\eta T$ and $\eta^{2}T^{2}$) can undermine merging performance. Thus, in the convergence phase, it is essential to control directional leakage (small $\delta$) and limit $\eta T$ to ensure high-quality merging. It supports previous empirical observations: fine-tuned models in current benchmarks typically remain within the same basin near the base model. For example, merging Qwen2.5-Math (Yang et al., 2024a) and Qwen2.5-Coder (Hui et al., 2024) yields poor performance. This is likely due to excessive post-training, which causes significant parameter drift. This insight suggests that poor merging results may not reflect algorithmic flaws, but rather issues with the fine-tuned models. When selecting models from Hugging Face, it’s helpful to choose fine-tuned models that are better suited for merging to minimize multi-task degradation.
@@
We examine the weight magnitude distribution of task vectors across our benchmark (see Fig.˜2 (a-b)). Our analysis reveals that InternVL2.5, which undergoes full fine-tuning, exhibits a right-skewed distribution. In contrast, Qwen2-VL, fine-tuned using LoRA, displays a multi-modal distribution. This is due to the low-rank nature and scaling factors of LoRA, which constrain the task vectors to be linear combinations in a reduced subspace, causing them to cluster along a few dominant magnitudes. Both models demonstrate distinct magnitude distribution patterns across different tasks. We also compute the normalized Frobenius norm of parameters (i.e., divided by the number of parameters). As shown in Fig.˜2 (c-d), the Frobenius norm varies significantly across tasks and layers, which presents challenges that we will address in our approach. The small task vector magnitudes suggest that fine-tuned models and base models exist in adjacent regions of the loss landscape with linear connectivity (Wu et al., 2023), facilitating effective model merging.
@@ 4.1 Merging Full Fine-tuning Models
By truncating singular values, we preserve critical features $V_{1:k}^{\top}$, which is similar to selecting principal components in Principal Components Analysis (PCA) (Abdi and Williams, 2010). This yields more accurate estimates of $\bm{x}_{i,l}$ than using $(\bm{\tau}_{i,l})^{\top}$.
@@ 5.1 Benchmark Details
VQA 588K GQA (en) (Hudson and Manning, 2019), VQAv2 (en) (Goyal et al., 2017), OKVQA (en) (Marino et al., 2019), LLaVA-Instruct (zh) (Liu et al., 2024a), CogVLM-Singleround (en&zh) (Wang et al., 2024c), CogVLM-Multiround (en&zh) (Wang et al., 2024c)
Geometry 190K GeoQA+ (zh) (Cao and Xiao, 2022), G-LLaVA (en) (Gao et al., 2023)
@@
Merging details. Following Task Arithmetic (Ilharco et al., 2023), we employ a single coefficient $\lambda$ to scale the merged vector before adding it to the base model. For all model merging methods, we determine the optimal merging coefficient $\lambda$ by searching within the range [0.1, 0.3, 0.5, 0.7, 1.0, 1.5]. In our implementation, the rank size $k$ in Eq.˜3 is simply defined as the rank of each task vector divided by the number of tasks (i.e., 5). We use the Adam optimizer with a learning rate of 1e-5 for InternVL while applying the SGD optimizer with a learning rate of 1e-4 for QwenVL. The number of optimization iterations is set to 300. We apply our method exclusively to the linear layer in the model.
@@ 5.2 Experimental Results
Table 2: Capability merging results on InternVL2.5 (full fine-tuning) across multiple tasks. For the merging methods, we highlight the best score in bold and the second-best score in underlining.
@@
Table 3: Capability merging results on Qwen2-VL (LoRA fine-tuning) across multiple tasks. For the merging methods, we highlight the best score in bold and the second-best score in underlining.
@@
Improved task vector optimization. Our method enhances task vector optimization stability, achieving optimal results. In Table˜4, we evaluate each component’s contribution to overall performance. Starting with WUDI Merging, we incrementally add one component at a time, reporting performance for both LoRA model merging (Qwen2-VL) and modality merging (Vicuna-7B). Replacing Adam with SGD alone does not necessarily improve performance; however, when combined with initializing the merged vector using the mean of task vectors, we observe a significant 4.43% improvement. Low-rank approximation further enhances performance, demonstrating its effectiveness in preserving critical knowledge from task vectors while maintaining the stability of the Frobenius norm. For full fine-tuned models, Tables˜2 and 6 show average improvements of 0.44% and 1.9% for OptMerge over WUDI Merging, respectively. This highlights the necessity of Eq.˜3 over Eq.˜1 for denoising task vectors and achieving robust merged-vector optimization.
@@
Rank size $k$. To further investigate the impact of rank size, we conduct additional ablation studies by setting $k$ to 10%, 20%, 30%, 40%, and 50% of the rank of each task vector. The results are summarized in Table 8. As shown, the performance remains relatively stable for $k$ ratios between 10% and 30%, indicating that OptMerge is robust to moderate changes in rank size.
Table 8: Ablation study on rank size $k$ ratio for OptMerge. The rank size $k$ is set to 10%, 20%, 30%, 40%, and 50% of the rank of each task vector.
$k$ ratio VQA Geometry Chart OCR Grounding Avg.
10% 30.90 57.26 51.49 18.42 68.40 76.10 46.39 76.36 69.99 73.96 56.93
20% 30.97 57.13 54.48 21.05 68.72 76.01 46.35 75.97 69.72 73.94 57.43
30% 31.55 57.15 54.50 21.05 68.72 76.27 45.67 73.63 66.84 70.92 56.63
40% 31.49 56.92 55.77 25.00 67.36 76.06 45.96 65.55 58.40 59.64 54.22
50% 31.37 56.68 56.75 23.68 68.08 75.81 45.02 61.45 54.80 56.19 52.98
Model scales. We extend our evaluation to the larger Qwen2.5-VL-32B-Instruct model and augment training with additional high-quality fine-tuning data. As shown in Table 9, OptMerge effectively combines multiple fine-tuned models while mitigating cross-task interference, achieving the best overall performance and surpassing the base Qwen2.5-VL-32B-Instruct. These results indicate that OptMerge remains effective and beneficial at larger model scales.
Table 9: Capability merging results on Qwen2.5-VL-32B-Instruct across multiple tasks. For the merging methods, we highlight the best score in bold.
Methods VQA Geometry Chart OCR Grounding Avg.
Qwen2.5-VL-32B-Instruct 41.39 59.34 79.21 47.36 83.64 79.62 64.58 88.01 82.41 84.06 70.96
Individual Geometry 42.67 60.25 80.34 43.42 86.76 80.83 66.54 89.58 83.72 84.56 71.87
Individual Grounding 41.60 59.61 78.86 42.10 85.88 79.65 64.19 88.32 82.95 84.04 70.72
Individual Chart 43.01 61.69 74.38 43.42 86.96 81.73 67.90 89.72 83.92 84.33 71.71
Individual VQA 42.24 62.75 78.40 42.11 86.68 81.04 67.06 89.74 83.90 84.72 71.86
Individual OCR 42.65 61.04 75.28 34.21 87.00 81.42 67.32 89.63 83.76 84.62 70.69
OptMerge (Ours) 43.52 62.50 80.01 43.42 88.92 81.91 66.37 89.94 83.97 84.68 72.52
Table 10: Evaluation of the merged model on general multimodal QA benchmarks.
MMMU DocVQA ScienceQA AI2D InfographicVQA
Individual Geometry 33.67 64.29 73.25 62.27 29.79
Individual Grounding 34.22 65.64 76.54 63.24 33.82
Individual Chart 30.33 57.13 40.01 29.86 26.02
Individual VQA 26.00 62.93 50.83 44.59 39.07
Individual OCR 38.00 77.67 63.66 54.39 41.97
OptMerge (Ours) 39.33 84.18 91.89 79.44 56.84
General tasks. We further evaluate the merged model (based on InternVL2.5-VL-1B) on a set of general multimodal QA benchmarks that require combinations of multiple abilities. The results are shown in Table 10. On these integrated benchmarks that require multiple abilities, single-ability models cannot solve the tasks effectively. In contrast, our OptMerge, which merges all specialized models, demonstrates emergent integrated capabilities and consistently outperforms the best individual model for each task, with an average improvement of 10.85% across benchmarks.